# Laypeople’s Egocentric Perceptions of Copyright for AI-Generated Art

### v1

@@ 2.5. Lay Perceptions of Copyright
Copyright protection exists to achieve several objectives (Lee et al., 2023). For instance, it promotes fairness by granting authors the right to exclusively control the fruits of their own labor (Bair, 2017). It also safeguards moral rights of creators, protecting the emotional bond between authors and their works (Damich, 1988). Most relevant to this research is copyright law’s aim to incentivize creativity. Copyright law attempts to promote creativity by granting exclusive rights over creative works to authors. These exclusive rights determine that only authors can profit from their creations, incentivizing them to continue exercising their creativity for further financial benefit (Mandel, 2014, 2016). Unless potential authors and rights-holders (i.e., laypeople) understand and agree with what is eligible for copyright protection and what rights are associated with it, copyright law may fail to incentivize the production of creative outputs. Hence, examining public perceptions and expectations of IP law is essential to ensure that current regulations effectively meet their goals and to identify potential changes if necessary.
Prior work looking at perceptions of IP law and authorship demonstrate that lay perceptions may be in conflict with what the law proposes. Laypeople perceive IP law’s main objective as preventing plagiarism, although its main aim is more utilitarian by promoting creativity through exclusive rights (Mandel, 2016). In the internet, content creators also have mistaken beliefs about the copyright terms of the platforms they use (Fiesler et al., 2016). This paper investigates laypeople’s perceptions of copyright law concerning works generated with the assistance of a GenAI model.
@@ 2.6. Lay Perceptions of GenAI Outputs
Recent work has explored how people use and perceive GenAI models and their outputs. While some report that people have an inherent bias against AI-generated text and images (Gu and Li, 2022), particularly among those with stronger anthropocentric beliefs in order to “protect” human creativity (Millet et al., 2023), other studies have found that laypeople prefer human-created works because AI outputs are perceived to be of lower-quality (Bower and Steyvers, 2021; Köbis and Mossink, 2021). At the same time, many studies demonstrate that people are bad at distinguishing between human- and AI-generated images and text (Köbis and Mossink, 2021; Longoni et al., 2022), potentially because of flawed heuristics used to determine whether something is machine-generated (Jakesch et al., 2023).
Research has also explored perceptions of GenAI outputs in relation to their ownership and authorship, albeit not in direct relation to their potential copyright protection. Human creators assisted by GenAI are attributed more credit than creators working alongside another human (Jago and Carroll, 2023). An experiment found that perceptions of authorship of the AI model and related actors can be manipulated by how anthropomorphized the machine is (Epstein et al., 2020). Similarly, perceptions of ownership over AI-generated text also vary depending on writing style (Kadoma et al., 2023). Another study suggests that people may not be contrary to the idea of granting copyright rights to the AI model themselves (Lima et al., 2020). Prior work has also discovered that laypeople believe artists whose creations are being used to train GenAI should be compensated (Kastrenakes and Vincent, 2023). We build upon this prior work and investigate people’s opinions regarding GenAI under the lens of copyright law, providing implications to the development and regulation of GenAI models.444The question of copyright and authorship can also be framed as a question of responsibility. For instance, who is responsible for an AI-generated work and, thus, should have the corresponding rights? Although extensive literature has explored how laypeople attribute responsibility for harms caused by AI (Lima et al., 2021a, b, 2023; Lee, 2021; Malle et al., 2015; Kim and Hinds, 2006), it is still an open question whether the results would replicate in the case of positive responsibility (i.e., credit for AI-generated works) (Porsdam Mann et al., 2023; Danaher and Nyholm, 2021).
@@ 3. Methods
We conducted a large-scale experiment to capture laypeople’s opinions concerning the potential copyright protection of AI-generated art. Furthermore, we studied whether people exhibit egocentric biases in their opinions by looking at the differences between the perceptions of those who use AI models to create art and those who observe and evaluate AI-generated art. In this section, we describe our experimental design, including the experimental setting and the procedures we employ to gather our data. We also formulate hypotheses exploring potential egocentric biases.
@@ 3.1. Setting
To study lay perceptions of AI-generated art, we held an online AI art exhibition. We recruited laypeple to participate in the exhibition as either creators or evaluators (or both, as explained below). Participants were told that the exhibition was juried, meaning that not all images would be displayed in the exhibition, and that an online website would display the best 10 images, the creators of which would be awarded a monetary award. Our setting not only enabled us to capture lay perceptions of AI-generated art in relation to copyright law more generally, but also allowed us to form hypotheses regarding how its incentive structure may influence these perceptions.
Our choice of experimental setting was motivated by a series of considerations. Our first consideration is related to the experiment’s ecological validity. We mimicked a real-world setting common in the art world: a juried art exhibition in which participants can submit their art online, and where the best rated submissions (i) receive recognition by being displayed in the exhibition and (ii) receive monetary rewards (see (Roose, 2022) for an equivalent scenario involving AI-generated art). The monetary rewards and recognition also serve a second purpose—incentivizing creators to put effort in creating images, as real-world creators of (GenAI or human-created) images normally would. Furthermore, the promise of monetary rewards and recognition to creators may also incentivize evaluators to take the task more seriously, since they are made aware of the fact that their responses influence other participants’ outcomes.
It is important to note that our exhibition differs from many juried art exhibitions in one important aspect: a part of the jury consists of creators’ peers who can also submit their art, and not an (unbiased) set of professionals who do not compete in the exhibition. We opted for this design choice to study egocentric effects in people’s perceptions of GenAI images and their potential copyright.
@@ 3.2. Experimental Conditions
Before providing more details about each of the three experimental conditions to participants, we introduced the study setting, which was similar across all three treatments. On the study’s landing page, we described our experiment and gathered participants’ informed consent. We explained the study setting, and—depending on the experimental condition to which they were assigned—informed participants that they will be generating and/or evaluating images for an AI-generated art exhibition. After reading the task description, participants were asked comprehension check questions to ensure that they understood the task; participants could not continue with the study until they answered the questions correctly. Furthermore, we informed participants that the creators of the 10 highest-rated submissions to our juried art exhibition would have their submissions displayed in an online gallery and earn a monetary reward of 25 USD. Finally, we provided participants with the following definition of GenAI models:
@@ 3.2.1. Creators
The experiment commenced with a three-step tutorial, in which participants were taught how to use the GenAI model to generate images. In the first step, participants were taught to write prompts. Participants were told that the AI model would generate an image according to their written instructions and were asked to instruct the AI model to generate “an image of a cat in a comic-book style of art”. In the second step, participants were informed that they could ask the GenAI model to generate as many images as they wanted. We emphasized that AI model they were using (DALL-E 3 (Betker et al., 2023)) did not keep previous instructions in its memory, meaning that they had to fully describe the image they intended to generate each time. As an exercise, participants were told to instruct the AI model to generate “an image of a rabbit in an abstract style of art”. Third, participants were shown how to navigate between all the images they created (with arrows located beside the image), and how to select the image they would like to submit for the consideration at the exhibition (by clicking on a button).
After completing the tutorial, creators were asked to generate their submission to our AI art exhibition (see Figure 2(a)). They were randomly assigned the task of generating a portrait, an image of a landscape, or an abstract image on a between-subjects basis to provide participants with some guidance and ensure some variance in our data. Participants could generate as many images as they wanted and select whichever image they preferred, independently of the order in which they were generated. After selecting the image they wanted to submit to the AI art exhibition, participants evaluated the image they submitted with respect to several variables, all of which we explain in Section 3.3.
@@ 3.2.2. Invested Evaluators
Similarly to creators, invested evaluators also completed a tutorial before generating, selecting, and evaluating their own submission to the AI art exhibition. After evaluating their own submission, they evaluated submissions made by four randomly selected creators, one at a time (see Figure 2(b)). When evaluating a creator’s submission, participants had access not only to the image the creator submitted to the AI art exhibition, but also to all images generated by that creator and their respective prompts, as well as some descriptive statistics about the generation process (namely, the number of images that the creator generated before deciding on their final submission and how long they took to create and select it).
@@ 3.2.3. Uninvested Evaluators
Unlike invested evaluators, uninvested evaluators did not have the chance to complete the tutorial, nor to create and submit an image to the AI art exhibition. Instead, participants only evaluated submissions made by four randomly selected creators, one at a time. They used the same interface as invested evaluators (see Figure 2(b)).
@@ 3.3. Measures
All participants evaluated their own creations and/or images created by others. Each image was judged with respect to four groups of questions. Each of the three first question groups addressed a separate research question (see Section 1), whereas the last group helped determine which art would be included in the AI art exhibition.
Factors associated with copyright decisions (RQ1):
@@
Effort: Although some jurisdictions have rejected that mere effort warrants copyright protection (i.e., the sweat of the brow doctrine), other countries, such as Australia, suggest that effort could be sufficient (Eshraghian, 2020). Scholars have also defended that copyright decisions regarding GenAI outputs could consider the effort put by users (Guadamuz, 2017). Perceived effort was evaluated by agreeing with the following statement on the same 7-point scale: “[I/The person] put in effort to generate the final image.”
Skills: There exists a legal precedent in Canada (200, 2004) that states that copyright protection requires exercise of non-trivial skills. Similarly, Australian law (196, 1964; Eshraghian, 2020) suggests that skill is sufficient for copyright. Participants evaluated skills by agreeing with the following statement on the same 7-point scale: “[I/The person] used [my/their] skills to generate the final image.”
Attribution of authorship (RQ2): The question of copyright is closely related to the question of authorship. One of the reasons why AI-generated art is not eligible for copyright protection in the US is that it lacks human authorship (Davis, 2023). Hence, deciding who is the author (or authors) of a GenAI output is crucial to determining whether it is eligible for copyright protection. Although much of the discussion surrounds authorship of the user and the AI model itself, it is also possible that laypeople perceive other actors as authors, such as the artists whose creations were used to train the GenAI model and the company that developed it. Hence, we asked participants to what extent they agreed that these four entities are authors using the same 7-point agreement scale:
@@
Attribution of rights (RQ3): Copyright law grants several exclusive rights to the holder, such as the right to distribute, reproduce, and display the work, as well as make copies and prepare derivative materials (usc, 2022). We captured participants opinions about two of these rights: 1) the right to display and 2) the right to make copies.
@@
Score evaluation: Participants also evaluated each image using a 11-point scale: “On a scale from 0 (Very bad) to 10 (Very good), how would you evaluate this image?” Responses to these questions determined which images were selected for the exhibition and thus received the monetary award.
@@
First, participants answered the groups of questions about factors associated with copyright decisions (1) and authorship (2). The order of these two groups of questions was randomized, as was the order of the questions within each group.555We did not provide a definition of what creativity, effort, and skills mean in the context of our study. We did so because of the lack of clear legal definition of these factors, as well as disagreement on whether these factors matter when deciding whether a work warrants copyright (see Section 2.2). Whether a work satisfies these conditions is open to legal interpretation and argument. Hence, we decided to not bias participants’ responses with respect to a particular definition of these factors and instead capture lay interpretations of these factors. Questions related to attribution of rights (3) followed, also in random order. Finally, participants scored the images to determine awards (4). For readability, some questions were rephrased depending on whether respondents were evaluating their own or other people’s creations (see above for the exact phrasing). Before completing the study, participants also answered a series of exploratory questions, such as how often they use image and text GenAI models and demographic questions.
3.4. Hypotheses
Particularly relevant for our study, people exhibit egocentric biases in their perceptions of creative works. Authors of creative works overestimate the value of their products (Buccafusco and Sprigman, 2010). Similarly, people value their self-made products as much as experts’ creations and expect others to share this view (Norton et al., 2012). Hence, people may value AI-generated art they create more than they value others’ art and more than others value their art. Moreover, people overestimate their own contributions to group projects (Ross and Sicoly, 1979) and take undue credit for achievements while attributing failures to external factors (Shepperd et al., 2008; Miller and Ross, 1975). Therefore, people may attribute higher degrees of authorship and more rights to themselves. Finally, creators overestimate the creativity of their own creations (Buccafusco and Sprigman, 2011)—a finding that we expect to replicate in our study. Combined, this body of research motivates us to study egocentric effects on people’s perceptions of copyright of AI-generated art.

### v2

@@ 2.3. Lay Perceptions of Copyright
Copyright protection exists to achieve several objectives (Lee et al., 2023). For instance, it promotes fairness by granting authors the right to exclusively control the fruits of their own labor (Bair, 2017). It also safeguards the moral rights of creators, protecting the emotional bond between authors and their works (Damich, 1988). Most relevant to this research is copyright law’s aim to incentivize creativity. Copyright law attempts to promote creativity by granting exclusive rights over creative works to authors. These exclusive rights determine that only authors can profit from their creations, incentivizing them to continue exercising their creativity for further financial benefit (Mandel, 2014, 2016). Unless potential authors and rights-holders (i.e., laypeople) understand and agree with what is eligible for copyright protection and what rights are associated with it, copyright law may fail to incentivize the production of creative outputs. Hence, examining public perceptions and expectations of IP law is essential to ensure that current regulations effectively meet their goals and to identify potential changes if necessary.
Prior work looking at perceptions of IP law and authorship demonstrates that lay perceptions may be in conflict with what the law proposes. Laypeople perceive IP law’s main objective as preventing plagiarism, although its main aim is more utilitarian by promoting creativity through exclusive rights (Mandel, 2016). In the internet, content creators also have mistaken beliefs about the copyright terms of the platforms they use (Fiesler et al., 2016). This paper investigates laypeople’s perceptions of visual art generated with the assistance of a GenAI model through the lens of copyright law.
@@ 2.4. Lay Perceptions of GenAI Outputs
Recent work has explored how people use and perceive GenAI models and their outputs. While some report that people have an inherent bias against AI-generated text and images (Gu and Li, 2022), particularly among those with stronger anthropocentric beliefs in order to “protect” human creativity (Millet et al., 2023), other studies have found that laypeople prefer human-created works because AI outputs are perceived to be of lower-quality (Bower and Steyvers, 2021; Köbis and Mossink, 2021). At the same time, many studies demonstrate that people are bad at distinguishing between human- and AI-generated images and text (Köbis and Mossink, 2021; Longoni et al., 2022; Mink et al., 2022), potentially because of flawed heuristics used to determine whether something is machine-generated (Jakesch et al., 2023; Mink et al., 2024).
Research has also explored perceptions of GenAI outputs in relation to their ownership and authorship, albeit not in direct relation to their potential copyright protection. Human creators assisted by GenAI are attributed more credit than creators working alongside another human (Jago and Carroll, 2023). An experiment found that perceptions of authorship of the AI model and related actors can be manipulated by how anthropomorphized the machine is (Epstein et al., 2020). Similarly, one’s perceived ownership over a piece of text they generated with AI assistance was found to vary depending on the AI’s writing style (Kadoma et al., 2023).
A few studies have also investigated people’s perceptions about issues related to copyright in the context of AI. A study that examined public perceptions about the possibility of granting AI various rights suggests that people may not be contrary to the idea of granting copyright rights to the AI model itself (Lima et al., 2020). A poll on people’s thoughts, feelings, and fears about AI found that laypeople believe artists whose creations are being used to train GenAI should be compensated (Kastrenakes and Vincent, 2023). We build upon this prior work to comprehensively investigate people’s opinions regarding GenAI through the lens of copyright law, providing implications to the development and regulation of GenAI models.
We note that the question of copyright and authorship can also be framed as a question of responsibility. For instance, who is responsible for an AI-generated work and, thus, should have the corresponding rights? Although extensive literature has explored how laypeople attribute responsibility for harms caused by AI (Lima et al., 2021a, b, 2023; Lee, 2021; Malle et al., 2015; Kim and Hinds, 2006), it is still an open question whether the results would replicate in the case of positive responsibility, such as attributing credit for AI-generated works (Porsdam Mann et al., 2023; Danaher and Nyholm, 2021), which we consider in this work.
2.5. Effects of Egocentricity, Competition, and Experience on Perceptions of Copyright
Particularly relevant for our study, people exhibit egocentric biases in their perceptions of creative works. Authors of creative works overestimate the value of their products (Buccafusco and Sprigman, 2010). Similarly, people value their self-made products as much as experts’ creations and expect others to share this view (Norton et al., 2012). Hence, people may value AI-generated art they create more than they value others’ art and more than others value their art. Moreover, individuals overestimate the creativity of their own creations (Buccafusco and Sprigman, 2011) and their own contributions to group projects (Ross and Sicoly, 1979), suggesting that similar egocentric biases may emerge in judgments of creativity and authorship, which help determine whether a creation is eligible for copyright protection (Hugenholtz and Quintais, 2021; 199, 1991). Combined, this body of research motivates us to study egocentric effects on people’s perceptions of copyright of AI-generated art.
Another bias that may emerge in the context of our study is what we call the competition effect. In our study, the highest-scored submissions received monetary rewards, meaning that participants’ outcomes depended on other people’s assessments. Some of those competing for the reward had the opportunity to inflate their own relative rating not only by giving high ratings to their own submission (in line with egocentric effects), but also by giving low scores to other people’s submissions. That is, they could benefit from sabotaging their competition. The problem of sabotage in competitions has been discussed extensively in economics: theory predicts its occurrence, and experiments show that sabotage is indeed empirically relevant (Chowdhury and Gürtler, 2015). Hence, creators of AI-generated art may sabotage others for their own financial interests. The competition effect is relevant in the context of copyright law because the financial interests it aims to protect might be susceptible to sabotage when competitors can determine whether something is eligible for copyright protection, as explored in our study. Therefore, we hypothesize that the incentive structure associated with our study may lead those with something at stake to judge others’ submissions more harshly than those who do not have anything to gain from doing so.
Finally, the experience of interacting with the GenAI model to create AI-generated art could impact one’s perceptions of outputs. Research has identified several psychological phenomena, such as the mere-exposure effect (Zajonc, 1968; Montoya et al., 2017), practice and learning effects (Donovan and Radosevich, 1999), and the curse of knowledge (Hinds, 1999), which lead us to hypothesize that participants’ evaluations may be influenced by their interaction (or lack thereof) with the GenAI model. Although all of these phenomena have their own idiosyncrasies, they all refer to effects due to experiences with or exposure to a piece of information or system. Hence, we also explore a potential experience effect, which posits that the experience of interacting with the GenAI model can influence participants’ perceptions about AI-generated art.
@@ 3. Methods
We conducted a large-scale experiment to capture laypeople’s opinions concerning copyright-relevant factors (the creativity, effort, and skill necessary to create art using AI), as well as the authorship and potential copyright protection of AI-generated art. Furthermore, we studied whether people exhibit biases in their opinions by evaluating hypotheses we form based on prior work regarding potential differences between the perceptions of those who use AI models to create art and those who observe and evaluate AI-generated art. In this section, we describe our experimental design, including the experimental setting and the procedures we employ to gather our data. The study was approved by our Ethics Review Board (ERB). All data and analysis scripts are available at https://tinyurl.com/GenAICopyrightOpinions.
@@ 3.1. Setting
To study lay perceptions of AI-generated art, we held an online AI art exhibition. We recruited laypeople to participate in the exhibition as either creators or evaluators (or both, as explained below). Participants were told that the exhibition was juried (Beyond the Veil Art, [n. d.]), meaning that not all images would be displayed in the exhibition, such that the art would be evaluated by a jury (in this case of other crowdworkers) and that an online website would display the 10 highest-rated images, the creators of which would be awarded a monetary award. Our setting not only enabled us to capture lay perceptions of AI-generated art in relation to copyright law more generally, but also allowed us to form hypotheses regarding how its incentive structure may influence these perceptions.
Our choice of experimental setting was motivated by a series of considerations. Our first consideration is related to the experiment’s ecological validity. We mimicked a real-world setting common in the art world: a juried art exhibition in which participants can submit their art online and where the best-rated submissions (i) receive recognition by being displayed in the exhibition and (ii) receive monetary rewards (see (Roose, 2022; Anderson, 2014) for an equivalent scenario involving AI-generated art). The monetary rewards and recognition also serve a second purpose—incentivizing creators to put effort into creating images, as real-world creators of (GenAI or human-created) images normally would. Furthermore, the promise of monetary rewards and recognition to creators may also incentivize evaluators to take the task more seriously since they are made aware of the fact that their responses influence other participants’ outcomes.
It is important to note that our exhibition differs from many juried art exhibitions in one important aspect: a part of the jury consists of creators’ peers who can also submit their art and not an (unbiased) set of professionals who do not compete in the exhibition. We opted for this design choice to study egocentric effects on people’s perceptions of GenAI images and their potential copyright.
@@ 3.2. Experimental Conditions
Before providing more details about each of the three experimental conditions to participants, we introduced the study setting, which was similar across all three treatments. On the study’s landing page, we described our experiment and gathered participants’ informed consent. We explained the study setting and—depending on the experimental condition to which they were assigned—informed participants that they would be generating and/or evaluating images for an AI-generated art exhibition. After reading the task description, participants were asked comprehension check questions to ensure that they understood the task; participants could not continue with the study until they answered the questions correctly. Furthermore, we informed participants that the creators of the 10 highest-rated submissions to our juried art exhibition would have their submissions displayed in an online gallery and earn a monetary reward of 25 USD. Finally, we provided participants with the following definition of GenAI models:
@@
The figure is divided into two subfigures. The left subfigure shows the study’s interface in which participants generated images using the AI model. The interface contains some instructions, a text box, buttons, and the generated image. The right sub-figure shows the interface used to evaluate the submitted images. The interface shows the generated images along with some metadata, followed by the survey questions.
@@ 3.2.1. Creators
The experiment commenced with a three-step tutorial, in which participants were taught how to use the GenAI model to generate images. In the first step, participants were taught to write prompts. Participants were told that the AI model would generate an image according to their written instructions and were asked to instruct the AI model to generate “an image of a cat in a comic-book style of art.” In the second step, participants were informed that they could ask the GenAI model to generate as many images as they wanted. We emphasized that AI model they were using (DALL-E 3 (Betker et al., 2023)) did not keep previous instructions in its memory, meaning that they had to fully describe the image they intended to generate each time. As an exercise, participants were told to instruct the AI model to generate “an image of a rabbit in an abstract style of art.” Third, participants were shown how to navigate between all the images they created (with arrows located beside the image) and how to select the image they would like to submit for consideration at the exhibition (by clicking on a button).
After completing the tutorial, creators were asked to generate their submission to our AI art exhibition (see left panel of Figure 2). They were randomly assigned the task of generating a portrait, an image of a landscape, or an abstract image on a between-subjects basis to provide participants with some guidance and ensure some variance in our data. Participants could generate as many images as they wanted and select whichever image they preferred, independently of the order in which they were generated. After selecting the image they wanted to submit to the AI art exhibition, participants evaluated the image they submitted with respect to several variables, all of which we explain in Section 3.3.
@@ 3.2.2. Invested Evaluators
Similarly to creators, invested evaluators also completed a tutorial before generating, selecting, and evaluating their own submission to the AI art exhibition. After evaluating their own submission, they evaluated submissions made by four randomly selected creators, one at a time (see right panel of Figure 2). When evaluating a creator’s submission, participants had access not only to the image the creator submitted to the AI art exhibition, but also to all images generated by that creator and their respective prompts, as well as some descriptive statistics about the generation process (namely, the number of images that the creator generated before deciding on their final submission and how long they took to create and select it).
@@ 3.2.3. Uninvested Evaluators
Unlike invested evaluators, uninvested evaluators did not have the chance to complete the tutorial nor to create and submit an image to the AI art exhibition. Instead, participants only evaluated submissions made by four randomly selected creators, one at a time. They used the same interface as invested evaluators (see right panel of Figure 2).
@@ 3.3. Measures
All participants evaluated their own creations and/or images created by others. Each image was judged according to four groups of questions. Each of the three first question groups addressed a separate research question (see Section 1), whereas the last group helped determine which art would be included in the AI art exhibition.
Factors associated with copyright decisions (RQ1):555We did not provide a definition of what creativity, effort, and skills mean in the context of our study because of the lack of clear legal definition of these factors, as well as disagreement on whether these factors matter when deciding whether a work warrants copyright (see Section 2.1). Whether a work satisfies these conditions is open to legal interpretation and argument. Hence, we decided not to bias participants’ responses with respect to a particular definition of these factors and instead capture lay interpretations of these factors. Had we provided a particular definition to participants, we would have biased them to answer our questions in a particular way. In Section 5, we explore whether lay opinions are aligned or at odds with existing legal definitions and interpretations.
@@
Effort: Although some jurisdictions have rejected that mere effort warrants copyright protection (i.e., the sweat of the brow doctrine in the US (199, 1991)), other countries, such as Australia, suggest that effort could be sufficient (Eshraghian, 2020). Scholars have also suggested that copyright decisions regarding GenAI outputs should consider the effort put in by users (Guadamuz, 2017). Perceived effort was evaluated by agreeing with the following statement on the same 7-point scale: “[I/The person] put in effort to generate the final image.”
Skills: There exists a legal precedent in Canada (200, 2004) that states that copyright protection requires the exercise of non-trivial skills. Similarly, Australian law (196, 1964; Eshraghian, 2020) suggests that skill is sufficient for copyright. Participants evaluated skills by agreeing with the following statement on the same 7-point scale: “[I/The person] used [my/their] skills to generate the final image.”
Attribution of authorship (RQ2): The question of copyright is closely related to the question of authorship. One of the reasons why AI-generated art is not eligible for copyright protection in the US is that it lacks human authorship (Davis, 2023). Hence, deciding who is the author (or authors) of a GenAI output is crucial to determining whether it is eligible for copyright protection. Although the user and the AI model itself could be seen as authors due to their role in the generation process, it is also possible that laypeople perceive other actors as authors, such as the artists whose creations were used to train the GenAI model and the company that developed it. Hence, we asked participants to what extent they agreed that these four entities are authors using the same 7-point agreement scale:
@@
Attribution of rights (RQ3): Copyright law grants several exclusive rights to the holder, such as the right to distribute, reproduce, and display the work, as well as make copies and prepare derivative materials (usc, 2022). We captured participants’ opinions about two of these rights: 1) the right to display and 2) the right to make copies.
@@
Score evaluation: Participants also evaluated each image using an 11-point scale: “On a scale from 0 (Very bad) to 10 (Very good), how would you evaluate this image?” Responses to these questions determined which images were selected for the exhibition and thus received the monetary award.
@@
First, participants answered the groups of questions about factors associated with copyright decisions (1) and authorship (2). The order of these two groups of questions was randomized, as was the order of the questions within each group. Questions related to attribution of rights (3) followed, also in random order. Finally, participants scored the images to determine awards (4). For readability, some questions were rephrased depending on whether respondents were evaluating their own or other people’s creations (see above for the exact phrasing). Before completing the study, participants also answered a series of exploratory questions, such as how often they use image and text GenAI models and demographic questions.
3.4. Hypotheses: Egocentricity, Competition, and Experience Effects